May 24, 2013 Installation

From GlueXWiki
Jump to: navigation, search

Meeting Time and Place

Friday, May 24, 2013 at 1:30 PM At Jefferson Lab, the meeting will be held in F326


To connect from the outside, please use ESNET

  1. ) ESNET: 8542553
  2. )To connect by telephone (supported on request only) dial:
  • You can look up toll-free number at
  • US and Canada: (866)740-1260 (toll free)
  • International: (303)248-0285 (toll call)
  • enter access code followed by the # sign: 3421244#

Up to GlueX/Hall-D Installation Meetings

Action Items From Last Meeting


  1. Announcements
  2. Updates to the Mechanical Installation Schedule ( Latest Rendition of Install Schedule) (Tim)
  3. Electronics Installation Schedule (Fernando) [| Schedule]
  4. Summer manpower
  5. Magnet Status
  6. Tagger Magnet Installation
  7. Microscope Installation
  8. Tagger Array Installation
  9. Pair Spectrometer Installation
  10. BCAL Installation
  11. FCAL Installation
  12. TOF Installation
  13. FDC Installation
  14. CDC Installation
  15. A.O.B.


Present: JLab (Tim, Elliott, Beni, Hovanes, Fernando, Mark, Lubomir); UConn (Jim); Regina (George); CMU (Curtis, Naomi); IU (Matt)

Mechanical Schedule: No impact if solenoid tests conclude in the next few weeks. Meeting on Tuesday to discuss a plan for going back to high current. By the end of next week we may have an idea how much time is required. If a refrigerator fix is needed that could take 6-8 weeks.

Electronics Schedule: No modifications to this schedule. Some concern about falling behind tagger microscope electronics and high-granularity pair spectrometer electronics. Review of cable trays to happen soon.

Summer Manpower: Students are required to take a safety orientation course SAF-099. Class must be taken before a badge is issued. Students are not allowed in the hall or on accelerator site alone for the first three months that their badge is issued.

Solenoid Mapping: along the center of the magnet was done before quench. Plan still includes going to a safe current and do mapping at that point.

Tagger Magnet: Steel and vacuum chamber have been installed and aligned. Vacuum chamber has been straightened and will be sealed soon to check O-rings between poles and chamber. Plan for mapping of the magnet in July.

Microscope: Specs for location, cooling, and other installation needs have been passed on to JLab.

Tagger Array: no report / Tim: still working on installation location and assembly structure.

Pair Spectrometer: no report / Tim: frame and adjustment for mechanical installation has been developed. Fernando: fiber readout has been settled and the electronics are a little behind.

BCAL: Surface preparation and new curing procedures appear to be working well. The plan is to not redo those light guides that have not shown any problems. Most problems have appeared about 24 hours after gluing. About 35-37% of BCAL readout modules have been completed and are ready to install.

FCAL: PMT base fabrication nearly complete. Dark room equipment on schedule to deliver around the second week of June. Cable trays will be in place for cabling in the summer

TOF: Installation plans are still in the early stages. Some frame parts may be constructed by the FSU shop.

FDC: Nothing new to report. Complete report made at last meeting. Mock-up cable installation is being done.

CDC: Plans have not changed. Still waiting for a drawing so that a high voltage cable support ring can be made.